Complete the ordered pair so that it is a solution to the equation
5x+y=3

Answer by ReadingBoosters(3246)

y = -5x + 3
when y=0
x = -5x=-3
x=3/5
when x=0
y = -5(0)+3
y=3
(3/5,3)
.................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.